# 1]Independent Researcher

### v1

@@ 2 PINNs for the (2+ 1)-dimensional VC-CNLSE
The PINNs methodology leverages the approximation capability of deep neural networks to solve differential equations through a minimization problem. A feed-forward neural network with an appropriate architecture (in terms of depth and width) is trained to satisfy both the governing equations and the boundary and/or initial conditions by minimizing a loss function that combines physics-based residuals and data-driven errors. This approach allows the network to learn complex relationships within the data, ensuring that the solution adheres to the underlying physical laws while also capturing any observed patterns in the data. In this section, we first introduce the general framework for PINNs methods applied to the (2+ 1)-dimensional coupled nonlinear equations with variable coefficients. Next, we introduce an enhanced PINN scheme and investigate the related modifications to the standard PINNs. In doing so, we consider the general (2+ 1)-dimensional variable coefficient coupled nonlinear equation in a complex environment space as follows:
@@
In the problem setting above, $\psi_{j}=\psi_{j}(\mathbb{X},t)$ represents a complex-valued function of the spatial variables $\mathbb{X}=(x,y)$ and the temporal variable $t$. $\mathcal{N}$ is a nonlinear operator acting on the solutions $\psi_{j}(\mathbb{X},t)$ for $j=1,2$, their derivatives with respect to the spatial variables, and is parametrized by $\lambda$, which can be an unknown parameter (constant or variable). We also denote the domain of spatial variables by $\Omega=[\mathbb{X}_{L},\mathbb{X}_{R}]$ and the corresponding domain boundary by $\partial\Omega$, where $\mathbb{X}_{L}=(x_{L},y_{L})$ and $\mathbb{X}_{R}=(x_{R},y_{R})$ represent the left and right boundaries, respectively. The computational interval for the temporal variable $t$ is also denoted by $[T_{0},T_{f}]$. For $j=1,2$, physics models based on the left-hand side of Eq. (4) can be formulated as:
@@
where $\mathcal{N}_{u_{1}}$, $\mathcal{N}_{v_{1}}$, $\mathcal{N}_{u_{2}}$, and $\mathcal{N}_{v_{2}}$ are specified nonlinear operators of the real-valued functions $u_{1}$, $v_{1}$, $u_{2}$, and $v_{2}$, respectively. A PINN model aims to predict these real-valued functions by minimizing a multi-objective loss function which is designed to minimize the discrepancy between the model’s predictions to the initial and boundary conditions (5) and (6), as well as the physical laws described by physics model in (8). The physics-informed part of the neural network model can be constructed using the provided physics models in (8), by differentiating the network’s outputs with respect to the spatial and temporal variables using AD technique. Considering a three-dimensional input data $\mathbb{z}^{0}=(x,y,t)$ to the neural network model, a PINN model comprises a fully connected neural network with depth $D$, $D-1$ hidden layers and an output layer. In each layer, we denote the number of neurons by $N^{\ell}$ for $\ell=1,\cdots,D$. By feeding the input data into the network, an affine transformation will be applied as follows:
@@
where $\mathbb{z}^{\ell-1}$ is the input data from the previous layer to the next layer. Before sending the input data to the next layer, a nonlinear activation function $\sigma^{\ell}(\cdot)$ should be applied component-wise to the transformed data of each layer. This activation function introduces nonlinearity into the model, allowing the neural network to capture complex patterns and relationships within the data. Common choices for the activation function include ReLU, sigmoid, or tanh, depending on the specific requirements of the problem being addressed [80]. In this paper, we use the tanh activation function for all our experiments. $\mathbb{w}^{\ell}\in\mathbb{R}^{N^{\ell}\times N^{\ell-1}}$ and $\mathbb{b}^{\ell}\in\mathbb{R}^{N^{\ell}}$ are the weight matrix and bias vector corresponding to the layer $\ell$, respectively. So, the final output of the neural network can be formed as
@@
where $\boldsymbol{\theta}:=\{(\mathbb{w}^{\ell},\mathbb{b}^{\ell})\}_{\ell=1}^{D}$ is the set of all trainable parameters, and the operator $\circ$ is the composition operator. The PINNs model aims to predict the solutions to the system (4)-(6) by optimizing the network’s trainable parameters. In doing so, we define the following mean square error (MSE) loss function, considering the residual forms of Eq. (5) and Eq. (6), along with the physics-informed component described by system (8)
@@
$Loss_{ic}(\boldsymbol{\theta})$ $:=\frac{1}{N_{ic}}\sum_{k=1}^{N_{ic}}w_{ic}^{(k)}\Big{(}\Big{\lvert}u_{1}^{k}-\hat{u}_{1,\boldsymbol{\theta}}(\mathbb{X}^{k},t=T_{0})\Big{\rvert}^{2}+\Big{\lvert}v_{1}^{k}-\hat{v}_{1,\boldsymbol{\theta}}(\mathbb{X}^{k},t=T_{0})\Big{\rvert}^{2}$
@@
$Loss_{bc}(\boldsymbol{\theta})$ $:=\frac{1}{N_{bc}}\sum_{k=1}^{N_{bc}}w_{bc}^{(k)}\Big{(}\Big{\lvert}u_{1}^{k}-\hat{u}_{1,\boldsymbol{\theta}}(\mathbb{X}^{k},t^{k})\Big{\rvert}^{2}+\Big{\lvert}v_{1}^{k}-\hat{v}_{1,\boldsymbol{\theta}}(\mathbb{X}^{k},t^{k})\Big{\rvert}^{2}$
@@
$Loss_{r}(\boldsymbol{\theta})$ $:=\frac{1}{N_{r}}\sum_{j=1}^{N_{r}}w_{r}^{(k)}\Big{(}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}$
$+\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}\Big{)},$
where $\Big{\{}(\mathbb{X}^{k},t=T_{0}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{ic}}$ and $\Big{\{}(\mathbb{X}^{k},t^{k}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{bc}}$ are the sets of training data corresponding to the initial and boundary conditions, respectively. We also denote the set of residual points by $\Big{\{}(\mathbb{X}^{k},t^{k})\Big{\}}_{k=1}^{N_{r}}$ for the residual networks $\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}$, $\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}$, $\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}$, and $\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}$. These points can be sampled either uniformly or randomly. $w_{ic}$, $w_{bc}$, and $w_{r}$ are point-specific weights that specify different weights assigned to the initial condition, boundary condition, and residual loss terms in the overall loss function. These weights allow for fine-tuning the relative importance of each component, enabling the model to balance the accuracy of the predictions across the different aspects of the problem. Then, the optimal values of the network’s parameters can be obtained by minimizing the MSE loss function (9) using variants of gradient descent algorithm (e.g., SGD and Adam) [81, 82], or other classes of optimization algorithms. The corresponding complex-valued physics-informed part of the PINN model for the (2+ 1)-dimensional VC-CNLSE (1) can be formulated as follows:
@@
where $n\geq 1$ is a pred-defined scalable factor. This configuration will introduce additional $\sum_{\ell=1}^{D-1}N^{\ell}$ trainable parameters to be optimized, where each neuron in each hidden layer has its own slope for the activation function. Then, the set of trainable parameters $\boldsymbol{\theta}$ consists of the set of weights and biases $\{(\mathbb{w}^{\ell},\mathbb{b}^{\ell})\}_{\ell=1}^{D}$ and the set of scalable parameters $\{a_{k}^{\ell}\}_{\ell=1}^{D-1}$, $\forall k=1,\cdots,N^{\ell}$. We initialize these scalable parameters such that $na_{k}^{\ell}=1,\,\,\,\forall n\geq 1$. In order to force the network to quickly increase the slope of the activation function, we define a slope recovery loss term to be added into the overall loss function as:
$Loss_{a}:=\frac{1}{\frac{1}{D-1}\sum_{\ell=1}^{D-1}\exp(\frac{\sum_{k=1}^{N^{\ell}}a_{k}^{\ell}}{N^{\ell}})}.$
@@
$Loss(\boldsymbol{\theta}):=Loss_{ic}(\boldsymbol{\theta})+Loss_{bc}(\boldsymbol{\theta})+Loss_{r}(\boldsymbol{\theta})+w_{a}\cdot Loss_{a}(\boldsymbol{\theta}).$ (11)
@@
In applying the proposed PINN framework for data-driven solution and parameter discovery to the model equation (1), the loss function (11) ensures that the model predictions satisfy the observed data corresponding to the initial and boundary conditions, as well as the physics-informed parts provided by the system of equations (10). In general, for a PINN model, the behavior of the equation residuals often exhibits many oscillations around zero. Indeed, the underlying PINN model aimes to force the equation residuals to zero by penalizing these residuals at so-called collocation points, which are sampled uniformly or randomly within the computational domain. The residual-based adaptive refinement (RAR) strategy is proposed to address this issue. The RAR strategy focuses on regions of the computational domain where the residuals are largest. This approach improves the distribution of residual points by adaptively sampling more points in these regions during the training phase. In all our experiments, we randomly sample m new residual points after specific training iterations until either the maximum number of iterations is reached or the value of the residuals falls below a predefined threshold. This motivates us to define region-specific weight $w_{r}^{reg}$ insteed of point-specific one for individual residual loss term in Eq. (11), corresponding to the physics-informed parts described by the governing equations.
For the system (1), we combine the proposed PINN framework with the RAR technique to demonstrate that each residual, as described in the system of equations (10), requires a different distribution of residual points. To evaluate the performance and accuracy of our method, we consider two metrics, the $\mathbb{L}^{2}$ error and the point-wise error, defined as follows:
@@
where $\psi_{j}(\cdot,\cdot)$ and $\hat{\psi}_{j,\boldsymbol{\theta}}(\cdot,\cdot)$ are the exact and predicted solutions, respectively, evaluated at sampled test points. The $\mathbb{L}^{2}$ error measures the overall deviation between the exact and predicted solutions across the entire domain, while the point-wise error assesses the deviation at specific points of interest. The architecture of the proposed PINN framework is depicted in Fig 1. Since we are working in a two-dimensional space, the input layer should be adjusted to include both spatial and temporal coordinates, which induces a three-dimensional problem solving. The input data is transformed by applying the weights and biases to each neuron in each hidden layer. An adaptive nonlinear activation function is then applied component-wise to the transformed data, activating each neuron. Each neuron in the previous layer is fully connected to the neurons in the next layer. In the output layer, there are four neurons, corresponding to the predictions for the system (1), which include the real and imaginary parts of the decomposed solutions. After the network’s outputs are obtained, the network is able to construct the physics-informed parts through the application of the AD technique. For this, the values of the variable coefficients must be computed first. Then, the physics-informed neural network can be constructed using the residuals provided by the physics models in (10). Both networks share their parameters during training. By constructing the loss functions corresponding to the governing equations, initial and boundary conditions, and the slope recovery loss, one can repeatedly obtain the optimal values of the network’s parameters by minimizing the overall loss function using appropriate optimization algorithms. In this paper, the loss functions are optimized using both the Adaptive Moment Estimation (Adam) and Limited-Memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithms. Adam is a gradient-based optimization algorithm that combines the advantages of two other extensions of stochastic gradient descent algorithm. It maintains adaptive learning rates for each parameter by computing first and second moments of the gradients. On the other hand, L-BFGS is a full-batch quasi-Newton method known for its effectiveness in optimizing functions by approximating the inverse Hessian matrix [83]. Combining these two methods leverages Adam’s robustness and adaptability in the initial stages of training with L-BFGS’s precision in fine-tuning the final solution. This minimization process will continue until the network’s error falls below a pre-defined threshold $\delta$ or the maximum number of iterations is reached. We initialize the network’s parameters using Xavier initialization technique.
@@
We sketch the proposed PINN algorithm for solving the (2+ 1)-dimensional VC-CNLSE in Algorithm 1. In all our experiments, we initialize the scalable parameters such that $n=5$ and $a_{k}^{\ell}=0.2$ for $\ell=1,\cdots,D-1$ and $k=1,\cdots,N^{\ell}$. We set the initial condition loss, boundary condition loss, and slope recovery loss weights as: $w_{ic}=1$, $w_{bc}=1$, and $w_{a}=\frac{1}{100}$. Setting $w_{a}=\frac{1}{100}$ ensures that the value of the loss is not too large. We generate the residual points randomly using Latin Hypercube Sampling (LHS) strategy [84] within the computational domain. We start training the constructed PINN model with an initial number of residual points, setting $m=25$. After $5000$ iterations, we add $2^{k}\times m$ ($k=1,2,3,...$) new residual points for each residual equations based on the RAR algorithm during the training phase. This approach helps to reduce the oscillations in the residual equations while using fewer residual points. The accompanying code for this paper is written in Python, and the numerical results are reported using the JAX library. The JAX library is used for its efficient automatic differentiation and high-performance numerical computing capabilities. The library’s Just-In-Time (JIT) compilation and automatic differentiation features significantly enhance the performance and accuracy of the computations involved in solving complex differential equations. The computations were performed on a Lenovo Legion laptop equipped with a 2.30 GHz Intel Core i7-11800H processor and a single NVIDIA GeForce RTX 3060 GPU card.
@@
Training data sets $\Big{\{}(\mathbb{X}^{k},T_{0}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{ic}}$ and $\Big{\{}(\mathbb{X}^{k},t^{k}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{bc}}$
@@
Construct a neural network model with Xiavier initialization technique
while  (error $<$ $\delta$ornIter$<$maxIterdo (a) Construct the physics-informed neural network by computing the variable coefficients and incorporating the network’s predictions into the physics models using the AD technique. (b) Specification of the overall loss function $Loss(\boldsymbol{\theta})$ using the losses corresponding to the governing equations, initial and boundary conditions, and the slope recovery loss function. (c) Add $m$ new residual points for each residual equation based on the RAR technique, where

### v2

@@ 2 PINNs for the (2+ 1)-dimensional VC-CNLSE
The PINNs methodology leverages the approximation capability of deep neural networks to solve differential equations through a minimization problem. A feed-forward neural network with an appropriate architecture (in terms of depth and width) is trained to satisfy both the governing equations and the boundary and/or initial conditions by minimizing a loss function that induces physics-informed components and data-driven solutions. This approach allows the network to learn complex relationships within the data, ensuring that the solution adheres to the underlying physical laws while also capturing any observed patterns in the data. In this section, we first introduce the general framework for PINNs methods applied to the (2+ 1)-dimensional coupled nonlinear equations with variable coefficients. Next, we introduce an enhanced PINN scheme and investigate the related modifications to the standard PINNs. In doing so, we consider the general (2+ 1)-dimensional variable coefficient coupled nonlinear equation in a complex environment space as follows:
@@
In the problem setting above, $\psi_{j}=\psi_{j}(\mathbb{x},t)$ represents a complex-valued function of the spatial variables $\mathbb{x}=(x,y)$ and the temporal variable $t$. $\mathcal{N}$ is a nonlinear operator acting on the solutions $\psi_{j}(\mathbb{x},t)$ for $j=1,2$, their derivatives with respect to the spatial variables, and is parametrized by $\lambda$, which can be an unknown parameter (constant or variable). We also denote the domain of spatial variables by $\Omega=(\mathbb{x}_{L},\mathbb{x}_{R})$ and the corresponding domain boundary by $\partial\Omega$. The computational interval for the temporal variable $t$ is also denoted by $T_{0},T_{f}]$. For $j=1,2$, physics models based on the left-hand side of Eq. (2) can be formulated as:
@@
where $\mathcal{N}_{u_{1}}$, $\mathcal{N}_{v_{1}}$, $\mathcal{N}_{u_{2}}$, and $\mathcal{N}_{v_{2}}$ are specified nonlinear operators of the real-valued functions $u_{1}$, $v_{1}$, $u_{2}$, and $v_{2}$, respectively. A PINN model aims to predict these real-valued functions by minimizing a multi-objective loss function which is designed to minimize the discrepancy between the model’s predictions to the initial and boundary conditions (3) and (4), as well as the physical laws described by physics models in (6). The physics-informed part of the neural network model can be constructed using the provided physics models in (6), by differentiating the network’s outputs with respect to the spatial and temporal variables using the AD technique. Considering a three-dimensional input data $\mathbb{z}^{0}=(x,y,t)$ to the neural network model, a PINN model comprises a fully connected neural network with depth $D$, D1D-1 hidden layers, and an output layer. In each layer, we denote the number of neurons by $N^{\ell}$ for $\ell=1,\cdots,D$. By feeding the input data into the network, an affine transformation will be applied as follows:
@@
where $\mathbb{z}^{\ell-1}$ is the input data from the previous layer to the next layer. Before sending the input data to the next layer, a nonlinear activation function $\sigma^{\ell}(\cdot)$ should be applied component-wise to the transformed data of each layer. This activation function introduces nonlinearity into the model, allowing the neural network to capture complex patterns and relationships within the data. In this paper, we use the tanh activation function for all our experiments. $\mathbb{w}^{\ell}\in\mathbb{R}^{N^{\ell}\times N^{\ell-1}}$ and $\mathbb{b}^{\ell}\in\mathbb{R}^{N^{\ell}}$ are the weight matrix and bias vector corresponding to the layer $\ell$, respectively. So, the final output of the neural network can be formed as
@@
where $\boldsymbol{\theta}:=\{(\mathbb{w}^{\ell},\mathbb{b}^{\ell})\}_{\ell=1}^{D}$ is the set of all trainable parameters, and the operator $\circ$ is the composition operator. The PINNs model aims to predict the solutions to the system (2)-(4) by optimizing the network’s trainable parameters. In doing so, we define the following overall loss function based on the mean square error (MSE) losses, considering the residual forms of Eq. (3) and Eq. (4), along with the physics-informed component described by system (6)
@@
$Loss_{ic}(\boldsymbol{\theta})$ $:=\frac{1}{N_{ic}}\sum_{k=1}^{N_{ic}}w_{ic}^{k}\Big{(}\Big{\lvert}u_{1}^{k}-\hat{u}_{1,\boldsymbol{\theta}}(\mathbb{x}^{k},t=T_{0})\Big{\rvert}^{2}+\Big{\lvert}v_{1}^{k}-\hat{v}_{1,\boldsymbol{\theta}}(\mathbb{x}^{k},t=T_{0})\Big{\rvert}^{2}$
@@
$Loss_{bc}(\boldsymbol{\theta})$ $:=\frac{1}{N_{bc}}\sum_{k=1}^{N_{bc}}w_{bc}^{k}\Big{(}\Big{\lvert}u_{1}^{k}-\hat{u}_{1,\boldsymbol{\theta}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+\Big{\lvert}v_{1}^{k}-\hat{v}_{1,\boldsymbol{\theta}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}$
@@
$Loss_{r}(\boldsymbol{\theta})$ $:=\frac{1}{N_{r}}\sum_{k=1}^{N_{r}}w_{r}^{k}\Big{(}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}$
$+\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}\Big{)},$
where $\Big{\{}(\mathbb{x}^{k},t=T_{0}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{ic}}$ and $\Big{\{}(\mathbb{x}^{k},t^{k}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{bc}}$ are the sets of training data corresponding to the initial and boundary conditions, respectively. We also denote the set of residual points by $\Big{\{}(\mathbb{x}^{k},t^{k})\Big{\}}_{k=1}^{N_{r}}$ for encoding the physical laws $\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}$, $\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}$, $\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}$, and $\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}$ during training phase, constructed in (6). These points can be sampled either uniformly or randomly within the computational domain. $w_{ic}$, $w_{bc}$, and $w_{r}$ are point-wise loss weights that specify different weights assigned to the initial condition, boundary condition, and residual loss terms in the overall loss function, respectively. These weights allow for fine-tuning the relative importance of each component, enabling the model to balance the accuracy of the predictions across the different aspects of the problem. Then, the optimal values of the network’s parameters can be obtained by minimizing the overall loss function (7) using variants of the gradient descent algorithm (e.g., SGD and Adam) [52, 53], or other classes of optimization algorithms. The corresponding complex-valued physics-informed part of the PINN model for the (2+ 1)-dimensional VC-CNLSE (1) can be formulated as follows:
@@
where $n\geq 1$ is a pre-defined scalable factor. This configuration will introduce additional $\sum_{\ell=1}^{D-1}N^{\ell}$ trainable parameters to be optimized, where each neuron in each hidden layer has its own slope for the activation function. Then, the set of trainable parameters $\boldsymbol{\theta}$ consists of the set of weights and biases $\{(\mathbb{w}^{\ell},\mathbb{b}^{\ell})\}_{\ell=1}^{D}$ and the set of scalable parameters $\{a_{k}^{\ell}\}_{\ell=1}^{D-1}$, $\forall k=1,\cdots,N^{\ell}$. We initialize these scalable parameters such that $na_{k}^{\ell}=1,\,\,\,\forall n\geq 1$. In order to force the network to quickly increase the slope of the activation function, the authors in [50] defined a slope recovery loss term to be added into the overall loss function as:
$Loss_{a}(\boldsymbol{\theta}):=\frac{1}{\frac{1}{D-1}\sum_{\ell=1}^{D-1}\exp(\frac{\sum_{k=1}^{N^{\ell}}a_{k}^{\ell}}{N^{\ell}})}.$
@@
$Loss(\boldsymbol{\theta}):=Loss_{ic}(\boldsymbol{\theta})+Loss_{bc}(\boldsymbol{\theta})+Loss_{r}(\boldsymbol{\theta})+w_{a}\cdot Loss_{a}(\boldsymbol{\theta}),$ (9)
@@
In applying the proposed PINN framework for data-driven solution and parameter discovery to the model equation (1), the loss function (9) ensures that the model predictions satisfy the observed data corresponding to the initial and boundary conditions, as well as the physics-informed parts provided by the system of equations (8). In general, for a PINN model, the behavior of the equation residuals often exhibits many oscillations around zero. Indeed, the underlying PINN model aims to force the equation residuals to zero by penalizing these residuals at so-called residual points. The residual-based adaptive refinement (RAR) strategy is proposed to address this issue. The RAR strategy focuses on regions within the computational domain where the residuals are large. This approach improves the distribution of residual points by adaptively sampling more points in these regions during the training phase. In all our experiments, we randomly sample m new residual points after specific training iterations until either the maximum number of iterations is reached or the mean equation residuals fall below a pre-defined threshold. This motivates us to define region-specific weights for individual residual loss terms in Eq. (9), corresponding to the physics-informed parts described by the governing equations. Indeed, the loss term corresponding to the physics-informed parts can be reformulated as follows:
$Loss_{r}(\boldsymbol{\theta})$ $:=\frac{1}{N_{r}}\sum_{k=1}^{N_{r}}\Big{(}w_{r}^{\mathcal{F}_{\hat{u}}}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{F}_{\hat{v}}}\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}$
$+w_{r}^{\mathcal{G}_{\hat{u}}}\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{G}_{\hat{v}}}\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}\Big{)},$
where $w_{r}^{\mathcal{F}_{\hat{u}}},w_{r}^{\mathcal{F}_{\hat{v}}},w_{r}^{\mathcal{G}_{\hat{u}}}$, and $w_{r}^{\mathcal{G}_{\hat{v}}}$ are region-specific weights to specify different importance of each residual equation during training. Indeed, the value of these weights can be regarded as the sampled residual points within the computational domain using the RAR strategy during training. The proposed enhanced PINN framework introduces a region-specific sampling strategy for residual equations, allowing for different distributions of residual points tailored to each residual equation. In the standard PINN framework, a large and often uniformly distributed set of residual points is used throughout training. Since PINNs rely on automatic differentiation to compute derivatives of network outputs with respect to input variables, increasing the number of residual points significantly adds to the computational complexity. Consequently, training a standard PINN with many residual points can be computationally expensive and time-consuming. In contrast, the enhanced PINN’s targeted sampling strategy reduces the total number of residual points required, thereby lowering the training time. Moreover, by customizing the distribution of residual points for each residual equation, the enhanced PINN achieves improved convergence speed and overall training efficiency.
@@
The architecture of the proposed PINN framework is depicted in Fig. 1, including an input layer for spatial and temporal coordinates, inducing a three-dimensional problem solving. The input data is transformed by applying the weights and biases to each neuron in each hidden layer. An adaptive nonlinear activation function is then applied component-wise to the transformed data, activating each neuron. Each neuron in the previous layer is fully connected to the neurons in the next layer. In the output layer, there are four neurons, corresponding to the predictions of the system (1), which include the real and imaginary parts of the decomposed solutions. After the network’s outputs are obtained, the network is able to construct the physics-informed parts through the application of the AD technique. For this, the values of the variable coefficients must be computed first. Then, the physics-informed neural network can be constructed using the residuals provided by the physics models in (8). Both networks share their parameters during training. By constructing the loss functions corresponding to the governing equations, initial and boundary conditions, and the slope recovery loss, one can repeatedly obtain the optimal values of the network’s parameters by minimizing the overall loss function using appropriate optimization algorithms. In this paper, the loss functions are optimized using both the Adaptive Moment Estimation (Adam) and Limited-Memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithms. Adam is a gradient-based optimization algorithm that combines the advantages of two other extensions of the stochastic gradient descent algorithm. On the other hand, L-BFGS is a full-batch quasi-Newton method known for its effectiveness in optimizing functions by approximating the inverse Hessian matrix [54]. Combining these two methods leverages Adam’s robustness and adaptability in the initial stages of training with L-BFGS’s precision in fine-tuning the final solution. This minimization process will continue until the network’s error falls below a pre-defined threshold $\delta$ or the maximum number of iterations is reached.
@@
We sketch the proposed PINN algorithm for solving the (2+ 1)-dimensional VC-CNLSE in Algorithm 1.
@@
Training data sets $\Big{\{}(\mathbb{x}^{k},T_{0}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{ic}}$ and
$\Big{\{}(\mathbb{x}^{k},t^{k}),u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{bc}}$
@@
Construct a neural network model with Xavier initialization technique
while  (error >> $\delta$   and  nIter<<maxIter) do (a) Construct the physics-informed neural network by computing the variable coefficients and incorporating the network’s predictions into the physics models using the AD technique. (b) Specification of the overall loss function $Loss(\boldsymbol{\theta})$ using the losses corresponding to the governing equations, initial and boundary conditions, and the slope recovery loss function. (c) Add $m$ new residual points for each residual equation based on the RAR technique, where